OpenStudy (anonymous):

6.03] Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (texaschic101):

I would use elimination because the y's already cancel each other out. 2x + 5y = -3 -15x - 5y = -10 --------------add -13x = - 13 x = 1 now sub 1 into either of the original equations to find y 2x + 5y = -3 2(1) + 5y = -3 2 + 5y = -3 5y = -3 -2 5y = -5 y = -1 Now lets check it... -15x - 5y = -10 -15(1) - 5(-1) = -10 -15 + 5 = -10 -10 = -10 (correct) Your solution is (1,-1)